# Interpretable Visualizations of Data Spaces for Classification Problems

### v1

@@ II.2 Exposing decision boundaries in toxicity classifications
What makes a given molecule toxic? In this case study, we use PCovC to reveal a latent space that discriminates chemicals based on their neurotoxicity. More specifically, we leverage the dataset and classification procedures discussed in Wu et al. [33], which itself contains a processed subset of 68 bioassay endpoints for 7,660 molecules from the Tox21 dataset [34]. For this case study, we use a subset of this dataset, focusing only on the three assays that measure the potential for compounds to disrupt the acetylcholinesterase pathway. Since acetylcholinesterase is responsible for terminating synaptic transmission in the brain, these assays are considered as good indicators of neurotoxicity[35]. Within the dataset, these assays are labeled “tox21-ache-p1 (-1)”, “tox21-ache-p3 (-1)”, and “tox21-ache-p5 (-1)”. We will refer to molecules / data points considered “active antagonists” for a given assay as “toxic” and “inactive” as “non-toxic.”
@@
The chemistry and connectivity of these molecules are provided in the form of SMILES strings, from which we generate the RDKit fingerprint [36], which encodes molecular topology by mapping atom connectivity and local structure into a compact, sparse binary representation. We then use PCovC to create a map that delineates chemicals by the “tox21-ache-p3 (-1)” assay.
First, we compare the mapping results of PCovC as a function of our mixing parameter, $\alpha$, in Fig. 2. As expected, a PCA map of this dataset reflects little distinction between toxic and non-toxic molecules, where a subsequent analysis via logistic regression labels all data points as “non-toxic,” i.e., the majority population of the dataset. As $\alpha$ is lowered and the classification task bears more weight in Eq. 5, the separation between classes becomes clearer and our predictive power improves, even superceding that of logistic regression on the full-dimensional data. It is worth noting that the best performance, unlike PCovR, comes from a lower $\alpha$ value than 0.5 – this is likely due to the class imbalance within this dataset, and demonstrates the benefits of the tunability of PCovC.
For comparison, linear discriminant analysis (LDA, App. A.2) produces a map very similar to PCovC at $\alpha=0$, wherein only the labels inform the projector matrix. While this still performs better in subsequent classification tasks than PCA (and marginally worse than analogous PCovC maps), consider the implications of using a single-dimensional map to analyze the decisions made by a classification algorithm. This single-dimensional map can only reflect one dimension of chemical diversity, and all other informative dimensions are lost. Therefore, molecules that are similar in this first dimension, but otherwise dissimilar, cannot be distinguished in subsequent analysis. Furthermore, the orthogonal dimensions, such as the “y” axis in Fig. 2 also carry scientific relevance – these dimensions identify molecular characteristics that do not influence the classification of the data points.
More concretely, consider the analysis of the map in Fig. 3. To identify molecules along the decision boundary, we computed the distances between all “toxic” and “non-toxic” points in the first eight components of PCovC space and found the pairs closest to each other. Shown are eight such pairs, which demonstrate that the model has determined that toxicity can change with small hydrocarbon functionalization, swapping of functional groups, or fluorination. While similar analysis can be achieved through visual inspection (e.g via chemiscope.org[37]), we identified these molecules through simple mathematical operations on the PCovC latent space – enabling higher throughput identification of boundary data points, which can then be used to generate new candidate molecules for toxicity screening.
Figure 4: Maps of the Tox21 dataset considering multiple assays for classification. Here, the leftmost column corresponds to single-assay maps. The center column corresponds to two-assay maps, and the right, center plot corresponds to a three-assay map. Each plot is given for the $\alpha$ that corresponds to the best classification score for the corresponding assays. In the lower right, the PCA map is shown for comparison. LDA is not possible with multitask classification. Marker color corresponds to the values for the three assays, shown in the table in the upper right. Opacity corresponds to the train/test split, with training points shown with transparency.
@@
Of the three assays, Assay 1 is the most sensitive indicator of toxicity. In our postprocessed dataset, 488 of the 5,767 (8.46%) compounds are labeled as toxic by this assay, compared to 193 (3.35%) and 146 (2.53%) compounds for the Assays 2 and 3, respectively. This means that there are significantly many more data points considered “toxic” by Assay 1 that are “non-toxic” according to Assays 2 and 3. The PCovC maps for these singular assays and their combinations are shown in Fig. 4. Note the number of red points, indicating molecules only considered “toxic” by Assay 1, and black points, indicating molecules considered toxic by all assays. Maps generated from solely Assays 2 and 3 position these red points with the “non-toxic” white points. The clearest separation occurs in the map built on Assays 1 and 3, which delineates both decision boundaries clearly and discretely. One can imagine doing similar analysis to that of Fig. 3 to identify the molecules along these boundaries to determine the differentiation by the two assays.
III Discussion
@@
In this work, we demonstrated that PCovC can be a transformative tool for visualizing data spaces that involve classification problems. With the acetylcholinesterase assays from the Tox21 dataset, we center our analysis on molecules that neighbor the decision boundary, and extract from our model that minimal changes in molecular structure can delineate toxic molecules from nontoxic molecules. We observed that, compared to PCA and LDA, combining classification performance and information preservation in a single mapping strategy can unveil decision boundaries while still preserving similarity between data points, and can also lead to improvements in classification performance. In the computational chemical analysis, these aspects can lead to furthering our ability to identify and interpret structure-property relationships, however, we anticipate that this method will be impactful in any discipline where understanding feature importance is scientifically meaningful.
IV Methods
IV.1 PCovC
@@
Obtaining a PCovC map requires a few key steps. First, we must obtain $\mathbf{Z}$ for use in Eq. 8 through fitting a linear classifier with $\mathbf{X}$ and $\mathbf{Y}$. The exact choice of linear classification model is flexible; for the analysis in this work, logistic regression with L2 regularization was chosen for simplicity. The projector from the feature space to the latent space can be mathematically derived through following the procedure in Helfrecht et al. [28], but with the evidence matrix $\mathbf{Z}$ used instead of the linear regression approximation $\mathbf{\hat{Y}}$. When $n_{\mathbf{samples}}>n_{\mathbf{features}}$, one should obtain the projectors by performing an eigendecomposition of a modified Gram matrix $\mathbf{\tilde{K}}$ (Eq.  9).
$\mathbf{P}_{\mathbf{XT}}=\left(\alpha\mathbf{X}^{T}+(1-\alpha)\mathbf{W}\mathbf{Z}^{T}\right)\mathbf{U}_{\tilde{\mathbf{K}}}\mathbf{\Lambda}_{\tilde{\mathbf{K}}}^{-\frac{1}{2}}$ (9a)
$\mathbf{P}_{\mathbf{TX}}=\mathbf{\Lambda}_{\tilde{\mathbf{K}}}^{-\frac{1}{2}}\mathbf{U}_{\tilde{\mathbf{K}}}^{T}\mathbf{X}$ (9b)
@@
In Eq. 9, $\mathbf{U}_{\mathbf{\tilde{K}}}$ and $\mathbf{\Lambda}_{\mathbf{\tilde{K}}}$ are obtained through performing an eigendecomposition of $\mathbf{\tilde{K}}$. When $n_{\mathbf{samples}}<n_{\mathbf{features}}$, it is instead faster to use a modified covariance matrix $\mathbf{\tilde{C}}$ (Eq. 10).
$\mathbf{P}_{\mathbf{XT}}=\left(\mathbf{X}^{T}\mathbf{X}\right)^{-\frac{1}{2}}\mathbf{U}_{\mathbf{\tilde{C}}}^{T}\mathbf{\Lambda}_{\mathbf{\tilde{C}}}^{\frac{1}{2}}$ (10a)
$\mathbf{P}_{\mathbf{TX}}=\mathbf{\Lambda}_{\mathbf{\tilde{C}}}^{-\frac{1}{2}}\mathbf{U}_{\mathbf{\tilde{C}}}^{T}\left(\mathbf{X}^{T}\mathbf{X}\right)^{\frac{1}{2}}$ (10b)
@@
IV.2 Tox21 Preprocessing
This dataset, released in Wu et al. [33], contains Tox21 [34] bioassay data for 68 end points and 7,660 compounds, which are provided as SMILES strings. To better enable a visual analysis of the compounds, we separately convert the compounds into ASE [38] objects using RDKit [36]; 10 compounds were dropped from the dataset as they resulted in a “bad conformer id” error from RDKit. We then follow the preprocessing scheme in Wu et al. [33], in which all compounds other than those with an “inactive” result or the most common active results are not considered for model fitting. For the three acetylcholinesterase assays considered in this study, the number of compounds labeled as active antagonists outnumbered those that were labeled as active agonists by a wide margin. Thus, the active antagonists were kept and the active agonists were discarded. We use the same train-test split provided in the dataset itself; for the rigorous analysis of the “tox21-ache-p3 (-1)” assay, the result is a dataset of 6,357 compounds (5,807 train, 550 test). For the multitask PCovC maps, because we discard chemicals that are not inactive or active antagonists for all three assays, the dataset consists of 5,767 compounds (5,277 train, 490 test). As in the original paper, we then convert the SMILES strings of each of the compounds into RDKit fingerprints with default fingerprint size (2,048 bits).
Appendix A Existing Methods for Data Visualization and Explainability
A.1 Linear and Kernel Principal Components Analysis

### v2

@@ 3.1 What makes a molecule toxic? Exposing decision boundaries in toxicity classifications
In this case study, we use PCovC to reveal a latent space that discriminates chemicals based on their neurotoxicity. More specifically, we leverage the dataset and classification procedures discussed in [35], which itself contains a processed subset of 68 bioassay endpoints for 7,660 molecules from the Tox21 dataset [36]. We use a subset of this dataset, focusing only on the three assays that measure the potential for compounds to disrupt the acetylcholinesterase pathway. Since acetylcholinesterase is responsible for terminating synaptic transmission in the brain, these assays are considered as good indicators of neurotoxicity [37]. Within the dataset, these assays are labeled “tox21-ache-p1 (-1)”, “tox21-ache-p3 (-1)”, and “tox21-ache-p5 (-1)”. We will refer to molecules / data points considered “active antagonists” for a given assay as “toxic” and “inactive” as “non-toxic.” Following procedure from [35], we generate the RDKit fingerprint [38], which encodes molecular topology by mapping atom connectivity and local structure into a compact, sparse binary representation. We then use PCovC to create a map that delineates chemicals by the “tox21-ache-p3 (-1)” assay.
Figure 3: Effect of mixing parameter $\alpha$ on the 2D map and classification performance (on an out-of-sample dataset) for the “tox21-ache-p3 (-1)” assay. White and blue points indicate “non-toxic” and “toxic” molecules, respectively. Opacity denotes train (translucent)/ test (opaque) split. In each panel is a confusion matrix showing the accuracy of logistic regression on the testing data as they appear in the resulting map, where “TN” indicates the number of “true negatives”, “FP” indicates the number of “false positives”, and so on. For comparison, logistic regression on the full-dimensional data obtains a confusion matrix of TP=17, TN=505, FP=9, FN=19.
First, we compare the mapping results of PCovC as a function of our mixing parameter, $\alpha$, in Fig. 3. As expected, a PCA map of this dataset reflects little distinction between toxic and non-toxic molecules, where a subsequent analysis via logistic regression labels all data points as “non-toxic,” i.e., the majority population of the dataset. As $\alpha$ is lowered and the classification task bears more weight in Eq. (1), the activity cliff becomes clearer and our predictive power improves, even superceding that of logistic regression on the full-dimensional data. It is worth noting that the best performance, unlike PCovR, comes from a lower $\alpha$ value than 0.5 – this is likely due to the class imbalance within this dataset, and demonstrates the benefits of the tunability of PCovC.
For comparison, LDA produces a map very similar to PCovC at $\alpha=0$, wherein only the labels inform the projector matrix. This still performs better in subsequent classification tasks than PCA (and marginally worse than analogous PCovC maps). However, consider the implications of using a single-dimensional map to analyze the decisions made by a classification algorithm. This single-dimensional map can only reflect one dimension of chemical diversity, and all other informative dimensions are lost. Therefore, molecules that are similar in this first dimension, but otherwise dissimilar, cannot be distinguished. Furthermore, the orthogonal dimensions, such as the “y” axis in Fig. 3 also carry scientific relevance – these dimensions identify molecular characteristics that do not influence the classification of the data points.
@@
More concretely, consider the analysis of the map in Fig. 4. To identify molecules along the decision boundary, we computed the distances between all “toxic” and “non-toxic” points in the first eight components of PCovC space and found the pairs closest to each other. The eight pairs shown demonstrate that the model has determined that an activity cliff can appear due to small hydrocarbon functionalization, swapping of functional groups, or fluorination. While similar analysis can be achieved through visual inspection (e.g via chemiscope.org [39]), we identified these molecules through simple analysis of the PCovC latent space – enabling higher throughput identification of boundary data points, which can then be used to identify new candidates for toxicity screening.
Figure 5: Maps of the Tox21 dataset considering multiple assays for classification, two double-assay maps (left, center) and a triple-assay map (right). Each plot is given for the $\alpha$ that corresponds to the best classification accuracy for the corresponding assays. Marker color corresponds to the values for the three assays, shown in the table to the right. Opacity corresponds to the train/test split.
@@
Of the three assays, Assay 1 is the most sensitive indicator of toxicity. In our postprocessed dataset, 488 of the 5,767 (8.46%) compounds are labeled as toxic by this assay, compared to 193 (3.35%) and 146 (2.53%) compounds for the Assays 2 and 3, respectively. This means that there are significantly many more data points considered “toxic” by Assay 1 that are “non-toxic” according to Assays 2 and 3. The PCovC maps for these singular assays and their combinations are shown in Fig. 5. Note the number of red points, indicating molecules only considered “toxic” by Assay 1, and black points, indicating molecules considered toxic by all assays. A map generated from solely Assay 2 positions these red points with the “non-toxic” white points (as shown in the Supplementary Information). The clearest separation occurs in the map built on Assays 1 and 2, which delineates both decision boundaries clearly and discretely. One can imagine doing similar analysis to that of Fig. 4 to identify the molecules along these boundaries to determine the differentiation by the two assays.
3.2 Parsing multi-class datasets: spectral analysis of organosulfur bonding environments
One key limitation of using purely unsupervised models for QSPR is their tendency, by design, to map only the highest-variance or highest-information dimensions. However, many datasets contain multiple splits of interest corresponding to different properties. We apply PCovC to X-ray absorption near edge structure (XANES) spectra that are used to characterize the types of bonds that may occur in organosulfur compounts. We use the dataset published in [40], which contains 769 XANES spectra of organosulfur molecules, and consider three “levels” of property classification: sulfur oxidation state (-2, 0, +2), type of sulfur bond (sulfide, thiocarbonyl, thiol/mercaptan, sulfoxide, and sulfone), and whether the sulfur is the member of a conjugated system (aromatic) or not (aliphatic). Note that the classifications by oxidation state and bond type are multiclass, thus, in these cases LDA can be used to generate a multidimensional latent space.
Figure 6: Analysis of organosulfur dataset for different classification targets. Rows correspond to classification based on oxidation state (top), bond type (center), and aromaticity (bottom). (left) Underlying XANES spectra, colored by the class delineations. Heatmaps underneath the spectra correspond to correlations with the first two principal covariates. The main figure corresponds to the underlying spectra. (right) Maps according to PCA, LDA, and PCovC, colored by the class delineations. The PCA map will be the same for all three analyses. In the upper right, we indicate the accuracy of a logistic regression model on the resulting 2D map for the offset testing data. For PCovC, $\alpha$ is chosen to optimize test classification accuracy, indicated in each figure.
As can be seen in the first row of Fig. 6, it is straightforward to identify the the signals corresponding to different oxidation states, as the data separates well based on this set of classes. Because of this, a logistic regression model trained on the 2D PCA latent space yields perfect test accuracy, and LDA only misclassifies one test compound. Because PCA is well-suited to separate oxidation states, we can prioritize information retention in our PCovC mapping strategy. Thus, a larger $\alpha$ value of 0.85 is ideal to show the diversity of oxidation states. Note that this yields a map that is qualitatively similar to the PCA projection, with some minute differences.
Bond types are not well-separated in the PCA projection, given that the -2 oxidation state contains three different types. As shown in Fig. 6, the sulfides and the thiols are the most difficult to separate, leading to a loss in accuracy for both PCA (87%) and LDA (85%) . However, a logistic regression model trained on the PCovC latent space is still able to achieve a 92% bond type test classification accuracy. This demonstrates how incorporating classification loss into our mapping procedures “refocuses” our model; in purely unsupervised schema, having a wealth of data within one class versus another leads to this data dominating the mapping task. Because our sulfoxides and sulfones are easily recognizable, even a small amount of classification loss included in our mapping procedure (with $\alpha=0.95$) leads to significantly better separation of the thiocarbonyls, sulfides, and thiols.
The most difficult task is separating the compounds by aromaticity, as this is a property that cross-cuts the “higher-order” classes. As is suggested by the maps in Fig. 6, the principal components seem to be capturing the oxidation state and bond type of the organosulfur compounds. However, because neither of these properties has a direct relationship with bond aromaticity, a logistic regression model trained from the PCA latent space struggles (accuracy of 48%). However, PCovC and LDA are each able to find latent spaces in which the compounds are well delineated by their aromaticity, with accuracies of 78% and 73%, respectively.
So, why was PCovC able to consistently outperform LDA, an algorithm designed to find latent spaces that best delineate classes? The first step in computing any LDA projection involves finding and projecting data to the lowest dimensional subspace shared by the set of class means. In doing so, any directions of the data orthogonal to this subspace are lost, which can result in LDA dropping important intra-class scatter. This loss of information manifests itself via a hyperconcentration of points within their classes, which results in LDA overfitting the training data and struggling to correctly predict the bonding environments of out-of-sample chemistries. With PCovC, on the other hand, there is explicit regularization within the classification procedures, and we can decide how much inter- and intra-class variance we wish to retain in our map through tuning $\alpha$.
3.3 Understanding feature importance in non-linear classification of metals/nonmetals
Finally, what if each of our features are physically meaningful but we require non-linear models for effective classification? In this instance, it is unclear how these meaningful features contribute to the learned weights of a classification algorithm. However, with robust structure-property maps, we can still interpret these non-linear feature relationships. In this case study, we apply a Kernelized PCovC (KPCovC) on the dataset of 4,916 inorganic materials published in [41] to understand the relationships between elemental properties and metallic behavior. Each of these materials is represented as a set of 136 properties of the constituent elements, and the learnable target is whether the material is a metal or a non-metal. Following the procedure in [41], we define a non-metal and metal as materials that do or do not have an electronic band gap, and extract the evidence matrix from a trained support vector classification (SVC) model, where $\mathbf{Z}\equiv\mathbf{KP}_{KZ},$ where $\mathbf{K}$ is a similarity matrix determined by a radial basis function (RBF) kernel, using a Gaussian width $\gamma$ of 0.01.
Figure 7: Projections of inorganic materials found in [41] using Kernel PCA, KDA, and Kernel PCovC. Color and opacity denote metal (blue) versus non-metal (red) and training (transparent) versus testing (opaque) splits, respectively. Above each map is the distribution of points in each class.
With this kernel matrix, we compute the Kernel PCA, Kernel Discriminant Analysis (KDA) and KPCovC projections of the elemental property data, shown in Fig. 7. As was shown in [41], we also observe that applying KPCovC with a support vector machine classifier optimizes classification performance and best exposes the decision boundary; using KPCovC with $\alpha=0.1$, the accuracy to 90% compared to 88% and 65% with the analogous KDA and KPCA, respectively.
Figure 8: Correlations of materials characteristics with the non-linear KPCovC and KPCA maps. (left) Absolute value of Pearson correlation coefficients for different input features with the resulting maps from Fig. 7. Darker squares denote stronger correlation. (right) Recoloring of KPCovC (main) and KPCA (inset) by feature values, from low (red) to high (blue) on a linear scale.
With the KPCovC map in Fig. 7, we could again perform analysis by computing distances and identifying boundary data points in the KPCovC latent space similar to Fig. 4. However, given that the features have physical significance, we instead propose to view the map from the lens of the feature set itself. More specifically, we can quantitatively determine the discriminating features in the data by computing the Pearson correlation coefficient between each feature and the first and second principal covariates. In Fig. 8, we demonstrate such an analysis, while also computing correlations between other latent space dimensions. We find that features based on the electronegativity and ionization energy of the constituent elements, while highly correlated with the first principal component, have only a moderate to weak correlation with the first principal covariate. This indicates that, while electronegativity and ionization energy objectively demonstrate the greatest variance in the data, our model does not associate either property with changes in metallic behavior (upper right panels of Fig. 8). The same conclusion can be reached by viewing the correlation with the KDA 1D projection, but in the case of KDA, the information these features provide is mostly discarded on the map. On the other hand, the KPCovC map retains this orthogonal information (as shown by their strong correlations with the second principal covariate) while also highlighting that they largely do not contribute to the class outcome. Instead, as shown in the lower right panels of Fig. 8, features related to the number of p electrons and Mendeleev number, which are not strongly correlated with the first principal component, show strong correlation with the first principal covariate.
4 Conclusions
@@
In this work, we demonstrated that hybrid supervised-unsupervised mapping techniques, namely principal covariates classification, can be transformative tools for visualizing data spaces through the lens of classification problems. We observed that, compared to PCA and LDA, combining classification performance and information preservation in a single mapping strategy can unveil decision boundaries while still preserving similarity between data points, and can even lead to improvements in classification performance. In computational chemical analysis, these aspects can lead to furthering our ability to identify and interpret structure-property relationships.
Appendix A Detailed methods used to obtain results in the main text
A.1 Principal Covariates Classification
PCovC is implemented in the open-source package scikit-matter [42], a scikit-learn-affiliated package focused on methods born out of chemical sciences. This package follows the same API as scikit-learn [43], and thus obtaining a PCovC map follows the same fit and transform procedure similar to obtaining PCA in scikit-learn.
@@ A.1 Principal Covariates Classification
Behind the curtain, obtaining a PCovC map requires a few key steps. First, we must obtain $\mathbf{Z}$ for use in Eq. (2) through fitting a linear classifier with $\mathbf{X}$ and $\mathbf{Y}$. The exact choice of linear classification model is flexible. The projector from the feature space to the latent space can be mathematically derived through following the procedure in [28], but with the evidence matrix $\mathbf{Z}$ used instead of the linear regression approximation $\mathbf{\hat{Y}}$. When $n_{\mathbf{features}}>n_{\mathbf{samples}}$, one should obtain the projectors by performing an eigendecomposition of a modified Gram matrix $\mathbf{\tilde{K}}$ (Eq.  3).
$\mathbf{P}_{XT}=\left(\alpha\mathbf{X}^{T}+(1-\alpha)\mathbf{W}\mathbf{Z}^{T}\right)\mathbf{U}_{\tilde{\mathbf{K}}}\mathbf{\Lambda}_{\tilde{\mathbf{K}}}^{-\frac{1}{2}}$ (3a)
$\mathbf{P}_{TX}=\mathbf{\Lambda}_{\tilde{\mathbf{K}}}^{-\frac{1}{2}}\mathbf{U}_{\tilde{\mathbf{K}}}^{T}\mathbf{X}$ (3b)
@@
In Eq. (3), $\mathbf{U}_{\mathbf{\tilde{K}}}$ and $\mathbf{\Lambda}_{\mathbf{\tilde{K}}}$ are obtained through performing an eigendecomposition of $\mathbf{\tilde{K}}$, and $\mathbf{W}$ represents the set of weights from the linear classifier fit between $\mathbf{X}$ and $\mathbf{Y}$. When $n_{\text{features}}<n_{\text{samples}}$, it is instead faster to use a modified covariance matrix $\mathbf{\tilde{C}}$ (Eq. (4)).
$\mathbf{P}_{XT}=\left(\mathbf{X}^{T}\mathbf{X}\right)^{-\frac{1}{2}}\mathbf{U}_{\mathbf{\tilde{C}}}^{T}\mathbf{\Lambda}_{\mathbf{\tilde{C}}}^{\frac{1}{2}}$ (4a)
$\mathbf{P}_{TX}=\mathbf{\Lambda}_{\mathbf{\tilde{C}}}^{-\frac{1}{2}}\mathbf{U}_{\mathbf{\tilde{C}}}^{T}\left(\mathbf{X}^{T}\mathbf{X}\right)^{\frac{1}{2}}$ (4b)
@@
A.1.1 Inner Product of Evidence Tensors
Depending on the problem of interest, it may be necessary to perform PCovC with a dataset that involves multiclass, multilabel classification. In these instances, it will be necessary to compute $\mathbf{ZZ}^{T}$ with a 3D $\mathbf{Z}$. This can be done following Eq. (5).
$(\mathbf{ZZ}^{T})_{ij}=\sum_{c=1}^{n_{\text{classes}}}\sum_{l=1}^{n_{\text{labels}}}\mathbf{Z}[i,c,l]\cdot\mathbf{Z}[j,c,l]$ (5)
This approach can also be extended to problems in which each label corresponds to a different amount of classes, as $n_{\text{classes}}$ can be treated as a variable that changes across labels.
A.2 Datasets
The three datasets used in this work have been published in previous studies.
A.2.1 Tox21
This dataset, released in [35], contains Tox21 [36] bioassay data for 68 end points and 7,660 compounds, which are provided as SMILES strings. To better enable a visual analysis of the compounds, we separately convert the compounds into ASE [44] objects using RDKit [38]; 10 compounds were dropped from the dataset as they resulted in a “bad conformer id” error from RDKit. We then follow the preprocessing scheme in [35], in which all compounds other than those with an “inactive” result or the most common active results are not considered for model fitting. For the three acetylcholinesterase assays considered in this study, the number of compounds labeled as active antagonists outnumbered those that were labeled as active agonists by a wide margin. Thus, the active antagonists were kept and the active agonists were discarded. We use the same train-test split provided in the dataset itself; for the rigorous analysis of the “tox21-ache-p3 (-1)” assay, the result is a dataset of 6,357 compounds (5,807 train, 550 test). For the multitask PCovC maps, because we discard chemicals that are not inactive or active antagonists for all three assays, the dataset consists of 5,767 compounds (5,277 train, 490 test). As in the original paper, we then convert the SMILES strings of each of the compounds into RDKit fingerprints with default fingerprint size (2,048 bits).
A.2.2 Sulfur Spectra
In this case study, we analyze the dataset, released in [40], that contains 769 sulforganic compounds and their corresponding XANES spectra, each having been computed using time-dependent density functional theory. We use the same train-test split as was used in the original paper for each of the three classification tasks, which results in a training dataset of 717 spectra and a test dataset of 52 spectra (6.8%). We follow the same steps for preprocessing as was done in the original paper, in which we first divide each spectra by the highest intensity value in the training dataset, then mean-center each feature.
A.2.3 Metal/Nonmetal Classification
The dataset used for metal/nonmetal classification was published in [41]. It contains a set of 4,916 inorganic materials, each described by the average, difference, maximum, and minimum of 34 properties of its respective constituent elements. Following the procedure in [41], we perform an 80%-20% train-test split (3932 train, 984 test) and scale each feature in the data to have zero mean and unit variance.
Appendix B Theory underlying existing methods for data visualization and explainability
B.1 Linear and Kernel Principal Components Analysis